OpenStudy (anonymous):

The perimeter of a rectangular garden is 114 meters. The length is 6 meters longer than twice the width. Find the dimensions of the garden.

OpenStudy (3psilon):

We know that the equations for the perimeter of a rectangle is P=2l+2w . l being lenght and w being width

OpenStudy (3psilon):

So now we will substitute what we are given into that equation

OpenStudy (3psilon):

p=114 L=6+2w

OpenStudy (3psilon):

114=2(6+2w)+2w

OpenStudy (3psilon):

So with some simplification and combining like terms and a little algebra we come to 114=12+6w 102=6w w=6

OpenStudy (3psilon):

And since we know W we can easily find the length

OpenStudy (anonymous):

so it would be 114=2l+6?

OpenStudy (3psilon):

Close but in the perimeter equation it is p=2L+2w so it would be 114 =2L+2(6)